OpenStudy (anonymous):

The area under the curve of y=1/x from x=a to x=5 is approximately 0.916 where 1<=a<5. Using your calculator, find a.

OpenStudy (anonymous):

Hint: this means\[\int_a^5 \frac{dx}{x} \approx 0.916.\]What is the left hand side equal to?

OpenStudy (anonymous):

I understand that much but I'm not sure where to go from there.

OpenStudy (anonymous):

Hint 2: \[\int \frac{dx}{x} = \ln|x| + C.\]Can you take it from there?

OpenStudy (anonymous):

OH!!! gotcha. Thanks! :)

OpenStudy (anonymous):

Yep! Good luck.